# Synergising Human-like Responses and Machine Intelligence for Planning in Disaster Response

### v2

@@ Abstract
Index Terms:
Cognitive control, Optimization, Smart systems
@@ II-C Problem Formulation
$x_{t+\tau|t}=\Phi x_{t+\tau-1|t}+\Gamma u_{t+\tau|t}$ (3b)
$x_{t|t}=x_{t|t-1}$ (3c)
$x_{t+\tau|t}^{\mathbf{p}}\notin\triangle(C^{i}_{1:t})$ $\forall i$ (3d)
$\Psi_{t}=g\left(\Psi_{t-1},A_{t}\right)$ (3e)
$S_{t}=\underset{i\in\{1,2\}}{\arg\max}~{}\Psi_{t}(i)$ (3f)
$x_{t+\tau|t}\in\mathcal{X},~{}u_{t+\tau|t},\in\mathcal{U},~{}i\in\{1,..,N_{t}\}$ (3g)
@@
Constraints in Eqs. (3b) and (3c) pertain to the UAV’s dynamic constraints. Next, the constraint in Eq. (3d) ensures that the agent’s planned positional state $x_{t+\tau|t}^{\mathbf{p}}$ avoids residing inside the convex hull outlined by the firefront’s path $C^{i}_{1:t}$, where (.$\triangle$(.)△ ( . ) denotes the convex-hull operator, and $C^{i}_{1:t}$ is the volume occupied by the firefront up to time-step $t$ as provided by the disaster warning system.
@@ III-A System 1 - Human-like Fast/Heuristic Planning
3:Initialize Tree $Tr$ rooted at $\hat{x}_{t|t}$, $x_{c}\leftarrow\hat{x}_{t|t}$
@@
5:     $\tilde{x}\leftarrow$ $\mathcal{E}$)
6:     $x_{\text{nearest}}\leftarrow$ NearestNeighbour($Tr$, $\tilde{x}$)
@@
8:     if PathFree($x_{\text{nearest}},x_{\text{new}},\hat{C}_{t}$then
9:         Add $x_{\text{new}}$ to $Tr$
@@ III-B System 2 - Slow/Deliberative Optimization
$x_{t+\tau|t}=\Phi x_{t+\tau-1|t}+\Gamma u_{t+\tau|t}$ (5b)
$x_{t|t}=x_{t|t-1}$ (5c)
$\text{dot}(\alpha_{i,l},x^{\mathbf{p}}_{t+\tau|t})>\beta_{i,l}-Mb_{\tau,i,l}$ $\hskip 2.84526pt\forall\tau,i,l$ (5d)
$\sum_{l=1}^{6}b_{\tau,i,l}\leq 5$ $\hskip 2.84526pt\forall\tau,i$ (5e)
$b_{\tau,i,l}\in\{0,1\},~{}i\in\{1,..,\hat{N}_{t}\}$ $\hskip 2.84526pt\forall\tau,i,l$ (5f)
$x_{t+\tau|t}\in\mathcal{X},~{}u_{t+\tau|t}\in\mathcal{U}$ $\hskip 2.84526pt\forall\tau$ (5g)
@@
To accomplish this for all time-steps and all firefronts the constraint in Eq. (5d) uses the binary variable $b_{\tau,i,l}$ to check if Eq. (6) is violated, where $M$ is a big positive constant. Constraint (5e) counts the number of violations and ensures that is less or equal to $6$, thereby enforcing Eq. (6).
@@ III-C Supervisory Controller - System 1 and 2 Orchestration
The tuning hyperparameters $w_{ic}$ and $w_{id}$, for each $i\in\{1,..,3\}$, are set to establish the upper and lower limits of the attention parameters, respectively. In contrast, $w_{ia}$ and $w_{ib}$ adjust the attention response in relation to the input stimuli, and finally, the term $d_{t}=||x_{t}^{\mathbf{p}}-\hat{y}_{t}^{\mathbf{p}}||_{2}$ quantifies the Euclidean distance at time-step $t$ between the UAV’s current position, represented as $x_{t}^{\mathbf{p}}$, and the location of the nearest firefront, $\hat{y}_{t}^{\mathbf{p}}$, which emerged at time-step $t$. The dynamic behavior of $\Xi_{t}$ based on Eq. (9) for a specific set of hyperparameters is illustrated in Fig. 3.
@@ IV-A Simulation Setup
We evaluate the proposed approach by simulating the disaster scenario described in Sec. II according to the setup described below. For illustrative purposes, and in order to simplify the analysis we demonstrate our results in a top-down view (i.e., 2-dimensional settings). Therefore, the disaster environment $\mathcal{E}$ has dimensions $1$ km $\times$ $1$ km, and the UAV agent evolves (assuming a fixed altitude) according to the dynamics described in Eq. (1), with parameters $\Delta T$, $\zeta$, and $m$ are set to 1s, 0.2, and 1.05kg respectively. The UAV agent can reach a maximum velocity of 15m/s, utilizing an input control force in the range of $[-7.5,7.5]$N. The UAV’s initial position is set to $(x,y)=(50,50)$, and the goal region $G$ is represented as a rectangular region centered at $G_{o}=(900,900)$ with size 100m $\times$ 100m. The firefront process dynamics are given by $p_{b}=0.1$, $p_{s}=0.75$ and the birth density $b_{t}(.)$ is uniform inside the environment. In addition, $\bar{y}^{\nu}=3$m/s, $\bar{y}^{d}=\pi/4$, $\sigma_{\nu}=0.8$, $\sigma_{d}=\pi/20$, and finally $\gamma=0.8$. Moreover, the extend (i.e., size) of each firefront at each time-step is represented as a rectangle with dimensions uniformly sampled from the range $[100,250]$m. Subsequently, System 1 is configured for running for maximum of 500 iterations, with a step-size of 15m, and tracked with an LQR controller with $Q_{\text{LQR}}=2I_{4\times 4}$, and $R_{\text{LQR}}=0.1I_{2\times 2}$ Then the parameters in System 2’s objective function in Eq. (4) are set to $(\kappa_{1}=1,\kappa_{2}=0.001)$ if the attention is on the mission completion time attribute, and conversely set to $(\kappa_{1}=0.001,\kappa_{2}=1)$ if the attention is on the energy efficiency. Then, the parameter $M$ in Problem (P2) is set to $M=10^{5}$, and $T=80$. Regarding the supervisory controller, the performance matrix is set to $Q=\left[\begin{smallmatrix}0.8&0.5&0.3\\ 0.4&0.9&0.9\end{smallmatrix}\right]$, and the attention parameters $\Xi_{t}=\left(\xi_{1}(t),\xi_{2}(t),\xi_{3}(t)\right)$ take their values according to Eq. (9) with parameters $w_{1a}=40$, $w_{1b}=0.02$, $w_{1c}=9$, $w_{1d}=1$, $w_{2a}=10$, $w_{2b}=0.05$, $w_{2c}=7$, $w_{2d}=1$, $w_{3a}=10$, $w_{3b}=0.15$, $w_{3c}=7$, and $w_{3d}=1$, which generate the profiles illustrated in Fig. 3. Finally, the state transition matrix $H$ of the decision-making dynamics shown in Eq. (10) is set to $H=0.5I_{2\times 2}$. Finally, Problem (P2) is solved using the Gurobi solver, running on a 3.2GHz CPU system.
@@ IV-B Results
Figure 7: Comparison between the proposed tripartite Cognitive Controller (CC), and an agent that only utilized either System 1 (S1) or System 2 (S2).

### v3

@@ II-C Problem Formulation
$\hskip-8.53581ptx_{t+\tau|t}=\Phi x_{t+\tau-1|t}+\Gamma u_{t+\tau|t}$ (3b)
$\hskip-8.53581ptx_{t|t}=x_{t|t-1}$ (3c)
$\hskip-8.53581ptx_{t+\tau|t}^{\mathbf{p}}\notin\triangle(C^{i}_{1:t})$ $\hskip-28.45274pt\forall i$ (3d)
$\hskip-8.53581pt\Psi_{t}=g\left(\Psi_{t-1},A_{t}\right)$ (3e)
$\hskip-8.53581ptS_{t}=\underset{i\in\{1,2\}}{\arg\max}~\Psi_{t}(i)$ (3f)
$\hskip-8.53581ptx_{t+\tau|t}\in\mathcal{X},~u_{t+\tau|t},\in\mathcal{U},~i\in\{1,..,N_{t}\}$ (3g)
@@
Constraints in Eqs. (3b) and (3c) pertain to the UAV’s dynamic constraints. Next, the constraint in Eq. (3d) ensures that the agent’s planned positional state $x_{t+\tau|t}^{\mathbf{p}}$ avoids residing inside the convex hull outlined by the firefront’s path $C^{i}_{1:t}$, where $\triangle(.)$ denotes the convex-hull operator, and $C^{i}_{1:t}$ is the volume occupied by the firefront up to time-step $t$ as provided by the disaster warning system.
@@ III-A System 1 - Human-like Fast/Heuristic Planning
3:Initialize Tree TrTr rooted at $\hat{x}_{t|t}$, $x_{c}\leftarrow\hat{x}_{t|t}$
@@
5:  $\tilde{x}\leftarrow$ SampleRandomState($\mathcal{E}$)
6:  $x_{\text{nearest}}\leftarrow$ NearestNeighbour(TrTr, $\tilde{x}$)
@@
8:  if PathFree($x_{\text{nearest}},x_{\text{new}},\hat{C}_{t}$) then
9:   Add $x_{\text{new}}$ to TrTr
@@ III-B System 2 - Slow/Deliberative Optimization
$\hskip-8.53581ptx_{t+\tau|t}=\Phi x_{t+\tau-1|t}+\Gamma u_{t+\tau|t}$ (5b)
$\hskip-8.53581ptx_{t|t}=x_{t|t-1}$ (5c)
$\hskip-8.53581pt\text{dot}(\alpha_{i,l},x^{\mathbf{p}}_{t+\tau|t})\geq\beta_{i,l}+\epsilon-Mb_{\tau,i,l}$ $\hskip 2.84526pt\forall\tau,i,l$ (5d)
$\hskip-8.53581pt\sum_{l=1}^{6}b_{\tau,i,l}\leq 5$ $\hskip 2.84526pt\forall\tau,i$ (5e)
$\hskip-8.53581ptb_{\tau,i,l}\in\{0,1\},~i\in\{1,..,\hat{N}_{t}\}$ $\hskip 2.84526pt\forall\tau,i,l$ (5f)
$\hskip-8.53581ptx_{t+\tau|t}\in\mathcal{X},~u_{t+\tau|t}\in\mathcal{U}$ $\hskip 2.84526pt\forall\tau$ (5g)
@@
To accomplish this for all time-steps and all firefronts the constraint in Eq. (5d) uses the binary variable $b_{\tau,i,l}$ to check if Eq. (6) is violated, where $M$ is a big positive constant, and tolerance $\epsilon>0$. Constraint (5e) counts the number of violations and ensures that is less or equal to $6$, thereby enforcing Eq. (6).
@@ III-C Supervisory Controller - System 1 and 2 Orchestration
[…]
@@ IV-A Simulation Setup
We evaluate the proposed approach by simulating the disaster scenario described in Sec. II according to the setup described below. For illustrative purposes, and in order to simplify the analysis we demonstrate our results in a top-down view (i.e., 2-dimensional settings). Therefore, the disaster environment $\mathcal{E}$ has dimensions $1$ km $\times$ 11 km, and the UAV agent evolves (assuming a fixed altitude) according to the dynamics described in Eq. (1), with parameters $\Delta T$, $\zeta$, and $m$ are set to 1s, 0.2, and 1.05kg respectively. The UAV agent can reach a maximum velocity of 15m/s, utilizing an input control force in the range of [7.5,7.5][-7.5,7.5]N. The UAV’s initial position is set to $(x,y)=(50,50)$, and the goal region $G$ is represented as a rectangular region centered at $G_{o}=(900,900)$ with size 100m $\times$ 100m. The firefront process dynamics are given by $p_{b}=0.1$, $p_{s}=0.75$ and the birth density $b_{t}(.)$ is uniform inside the environment. In addition, $\bar{y}^{\nu}=3$m/s, $\bar{y}^{d}=\pi/4$, $\sigma_{\nu}=0.8$, $\sigma_{d}=\pi/20$, and finally $\gamma=0.8$. Moreover, the extend (i.e., size) of each firefront at each time-step is represented as a rectangle with dimensions uniformly sampled from the range [100,250][100,250]m. Subsequently, System 1 is configured for running for maximum of 500 iterations, with a step-size of 15m, and tracked with an LQR controller with $Q_{\text{LQR}}=2I_{4\times 4}$, and $R_{\text{LQR}}=0.1I_{2\times 2}$ Then the parameters in System 2’s objective function in Eq. (4) are set to $(\kappa_{1}=1,\kappa_{2}=0.001)$ if the attention is on the mission completion time attribute, and conversely set to $(\kappa_{1}=0.001,\kappa_{2}=1)$ if the attention is on the energy efficiency. Then, the parameter $M$ in Problem (P2) is set to $M=10^{5}$, and $T=80$. Regarding the supervisory controller, the performance matrix is set to $Q=\left[\begin{smallmatrix}0.8&0.5&0.3\\ 0.4&0.9&0.9\end{smallmatrix}\right]$, and the attention parameters $\Xi_{t}=\left(\xi_{1}(t),\xi_{2}(t),\xi_{3}(t)\right)$ take their values according to Eq. (9) with parameters $w_{1a}=40$, $w_{1b}=0.02$, $w_{1c}=9$, $w_{1d}=1$, $w_{2a}=10$, $w_{2b}=0.05$, $w_{2c}=7$, $w_{2d}=1$, $w_{3a}=10$, $w_{3b}=0.15$, $w_{3c}=7$, and $w_{3d}=1$, which generate the profiles illustrated in Fig. 3. Finally, the state transition matrix $H$ of the decision-making dynamics shown in Eq. (10) is set to $H=0.5I_{2\times 2}$. Finally, Problem (P2) is solved using the Gurobi solver, running on a 3.2GHz CPU system.
@@ IV-B Results
Figure 7: Comparison between the proposed tripartite Cognitive Controller ($C$), and an agent that only utilized either System 1 (S1) or System 2 (S2).